OpenStudy (anonymous):

how do i Determine the solution set of (x - 3)2 = 49?

OpenStudy (raden):

(x - 3)2 = 49 (x - 3)2 = 7^2 (x - 3)2 - 7^2 = 0 then factor out by using the difference square formula : a^2 - b^2 = (a+b)(a-b) therefore, (x - 3)2 - 7^2 = 0 can be written as (x - 3 + 7)(x - 3 - 7) = 0 (x + 4)(x - 10) = 0 now, solve for x